A cylindrical blood vessel is partially blocked by the buildup of plaque. At one point, the plaque decreases the diameter of the vessel by 72.0%. The blood approaching the blocked portion has speed ?0. Just as the blood enters the blocked portion of the vessel, what is its speed ?, expressed as a multiple of ?0?
Answer #1

Diameter of the blood vessel before the blocked portion = D0

Cross-sectional area of the blood vessel before the block portion = A0 = D02/4

Speed of the blood in the blood vessel before the blocked portion = V0

The plaque decreases the diameter of the vessel by 72%.

Diameter of the blood vessel in the blocked portion = D = 0.28D0

Cross-sectional area of the blood vessel in the blocked portion = A = D2/4

Speed of the blood in the blood vessel in the blocked portion = V

By continuity equation,

A0V0 = AV

(D02/4)V0 = (D2/4)V

D02V0 = D2V

D02V0 = (0.28D0)2V

V = 12.755V0

Speed of the blood as it enter the blocked portion of the vessel = 12.755V0